OpenStudy (anonymous):

the perimeter of a rectangle is given by the formula: p = 2(1+w) solve for w

OpenStudy (anonymous):

first, distribute. p=2+2w subtract 2 from each side. p-2=2w divide by 2. p-2/2=w

OpenStudy (anonymous):

my mistake, its not a 1 its an L